# bibliography.bib

### v3

@@ 8 Appendix
8.1 Proof of lemmas related to the Gibbs sampler for a the variance component model
Proof of lemma 4.1.
First note that the unnormalized joint distribution of $(\vec{\theta},V,W,\mu)$ is written as follows,
I want to find $\frac{1}{L}\sup_{(\vec{\theta},v,w,\mu)}\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$ where $L$ is a lower bound on
$\int_{\mathbb{R}^{I}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\times\mathbb{R}}g(\vec{\theta},v,w,\mu)d(\vec{\theta},V,W,\mu)$
Step 1: Find an upper bound on $\sup_{(\vec{\theta},v,w,\mu)}\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$
The ratio $\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$ is as follows for $C_{1}=\left(\frac{b_{1}}{b_{1}-1}\right)^{a_{1}}\left(\frac{b_{2}}{b_{2}-1}\right)^{a_{2}}$,
$\frac{g(\vec{\theta},V,W,\mu)}{f_{\nu}(\vec{\theta},V,W,\mu)}$
$=C_{1}e^{-1/V}e^{-1/W}\frac{\left(\prod_{i=1}^{I}\frac{1}{\sqrt{2\pi V}}e^{-\frac{(\theta_{i}-\mu)^{2}}{2V}}\right)\left(\prod_{i=1}^{I}\frac{1}{(2\pi W)^{J_{i}/2}}e^{-\frac{\sum_{j=1}^{J_{i}}Y_{ij}^{2}-2J_{i}\theta_{i}\bar{Y}_{i}+J_{i}\theta_{i}^{2}}{2W}}\right)}{\left(\prod_{i=1}^{I}\frac{1}{\sqrt{2\pi W/J_{i}}}e^{-\frac{J_{i}(\bar{Y}_{i}-\theta_{i})^{2}}{2W}}\right)}$
$\leq C_{1}e^{-1/V}e^{-1/W}\frac{1}{(2\pi V)^{I/2}}\frac{(2\pi W)^{I/2}/\prod_{i=1}^{I}J_{i}^{1/2}}{(2\pi W)^{\sum_{i=1}^{I}J_{i}/2}}\prod_{i=1}^{I}e^{-\frac{\sum_{j=1}^{J_{i}}Y_{ij}^{2}-2J_{i}\theta_{i}\bar{Y}_{i}+J_{i}\theta_{i}^{2}}{2W}+\frac{J_{i}\bar{Y}_{i}^{2}-2J_{i}\theta_{i}\bar{Y}_{i}+J_{i}\theta_{i}^{2}}{2W}}$
$=C_{1}e^{-1/V}e^{-1/W}\frac{1}{(2\pi V)^{I/2}}\frac{1}{(2\pi W)^{(\sum_{i=1}^{I}J_{i}-I)/2}}\frac{1}{(\prod_{i=1}^{I}J_{i})^{1/2}}\prod_{i=1}^{I}e^{\frac{J_{i}\bar{Y}_{i}^{2}-\sum_{j=1}^{J_{i}}Y_{ij}^{2}}{2W}}$
$\leq C_{1}\frac{1}{(2\pi)^{(\sum_{i=1}^{I}J_{i})/2}}\frac{1}{(\prod_{i=1}^{I}J_{i})^{1/2}}\left(e^{-1/V}\frac{1}{V^{I/2}}\right)\left(e^{-1/W}\frac{1}{W^{(\sum_{i=1}^{I}J_{i}-I)/2}}\right)$
The inequality is because $\sum_{j=1}^{J_{i}}Y_{ij}^{2}\geq J_{i}\bar{Y}_{i}^{2}$.
The supremum of $\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$ is bounded as follows for
$C=\left(\frac{b_{1}}{b_{1}-1}\right)^{a_{1}}\left(\frac{b_{2}}{b_{2}-1}\right)^{a_{2}}\frac{1}{(2\pi)^{(\sum_{i=1}^{I}J_{i})/2}}\frac{1}{(\prod_{i=1}^{I}J_{i})^{1/2}}$,
$\sup_{(\vec{\theta},v,w,\mu)}\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$ $\leq C\sup_{(v,w)}\left(e^{-1/v}\frac{1}{v^{I/2}}\right)\left(e^{-1/w}\frac{1}{w^{(\sum_{i=1}^{I}J_{i}-I)/2}}\right)$
$=C\sup_{(v,w)}\left(\frac{1}{v^{a_{v}+1}}e^{-\frac{1}{v}}\right)\left(\frac{1}{w^{a_{w}+1}}e^{-\frac{1}{w}}\right)$
Where $a_{v}=I/2-1$ and $a_{W}=(\sum_{i=1}^{I}J_{i}-I)/2-1$.
Since both of the functions for $V$ and $W$ are proportional to the inverse gamma density function, the supremum occurs when $V^{*}=\frac{1}{I/2-1+1}=\frac{2}{I}$ and $W^{*}=\frac{2}{\sum_{i=1}^{I}J_{i}-I}$. So that,
$\sup_{(\vec{\theta},v,w,\mu)}\frac{g(\vec{\theta},v,w,\mu)}{f_{\nu}(\vec{\theta},v,w,\mu)}$ $\leq C\left(\frac{1}{(V^{*})^{a_{v}+1}}e^{-\frac{1}{V^{*}}}\right)\left(\frac{1}{(W^{*})^{a_{w}+1}}e^{-\frac{1}{W^{*}}}\right)$
Step 2: Find a lower bound on $\int_{\mathbb{R}^{I}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\times\mathbb{R}}g(\vec{\theta},v,w,\mu)d(\vec{\theta},V,W,\mu)$
To find a lower bound over the integral we estimate the above integral over a closed set $B\subset\mathbb{R}^{I}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\times\mathbb{R}$. Since the integral can be extremely small, we integrate over the log of the un-normalised density and then take the exponent. That is by Jensen’s inequality,
$L\geq\exp\left[\int_{B}\log(g(\vec{\theta},v,w,\mu))d(\vec{\theta},V,W,\mu)\right]$
By Lemma 3.2, the inequality holds. ∎
@@ 8.1 Proof of lemmas related to the Gibbs sampler for a the variance component model
8.2 Proof of lemmas related to the Gibbs sampler for a James-Stein estimator
@@
$=\frac{\Gamma(\alpha+(q-1)/2)}{\Gamma(\alpha)}\frac{1}{(2\pi\beta)^{(q-1)/2}}$

### v4

@@ 8 Appendix
8.1 Convergence diagnostics
8.1.1 Real Data Example 4.1
Gelman-Rubin statistic Refer to caption
Traceplot for iterations $1\leq n\leq 10$ Refer to caption
Figure 7: Gelman-Rubin statistic and traceplot for Real Data Example 4.1
8.1.2 Numerical Example 5.1
Iteration V W $\mu$
10 1.064556 1.046249 1.062693
20 1.026606 1.021977 1.034042
30 1.017409 1.013104 1.021825
40 1.013610 1.011070 1.014756
50 1.010736 1.008834 1.011396
60 1.009296 1.007351 1.010150
Table 10: Gelman-Rubin statistic at various iterations for Real Data Example 5.1. The range of $\mu$ is -2 to 0.
Traceplot Refer to caption
Figure 8: Traceplot for $\mu$, $V$, and $W$ for Real Data Example 5.1
8.1.3 Real Data Example 5.1
$n$ $V$ $W$ $\mu$ $\theta_{1}$
50 1.186270 2.052273 2.984139 2.792903
100 1.095662 1.591515 1.597947 1.587649
150 1.083923 1.454056 1.241374 1.215334
200 1.062822 1.208700 1.086050 1.097018
250 1.054723 1.074812 1.038007 1.056482
Table 12: Gelman-Rubin statistic at various iterations for Real Data Example 5.1. The range of $V$ is $(0.5,3.5)$, $W$ is $1000,20000$, and $\mu,\theta_{1}$ is $(1400,1600)$. Note that because the range of initial values to calculate the Gelman-Rubin statistic is larger than the initial values in our bound, the convergence rates cannot be blindly compared. Rather, the Gelman-Rubin statistic indicates that the chains move relatively slowly.
Traceplot Refer to caption
Figure 9: Traceplot for $\mu$, $V$, and $W$ for Real Data Example 5.1
8.1.4 Real Data Example 6.1
Gelman-Rubin statistic Refer to caption
Figure 10: Gelman-Rubin statistic of $\sigma^{2}_{n}$.
Figure 11: Traceplot of 10 simulations of $\sigma^{2}_{n}$ where $\sigma^{2}_{0}=100$.
8.2 Proof of Theorem 3.1
Proof of Theorem 3.1.
Let $X_{n},Y_{n}$ be two copies of the Markov chain such that $X_{n}=k_{\Theta_{n}}(X_{0})$ and $Y_{n}=k_{\Theta^{\prime}_{n}}(Y_{0})$ (see Section 2 for more details on the notation). If $X_{\infty}\sim\pi$, then $k_{\Theta^{\prime}_{n}}(X_{\infty})\sim\pi$.
First we show that $E[d(X_{n},X_{\infty})^{r}]=E\left[d(X_{n},Y_{n})^{r}\frac{\mathrm{d}\pi(Y_{0})}{\mathrm{d}\nu(Y_{0})}\right]$ where $\mathrm{d}\pi(Y_{0})/\mathrm{d}\nu(Y_{0})=f_{\pi}(Y_{0})/f_{\nu}(Y_{0})$ and $\mathrm{d}\pi/\mathrm{d}\nu$ is the Radon-Nikodym derivative.
$E[d(X_{n},X_{\infty})^{r}]$ (15)
$=E[E[d(k_{\Theta_{n}}(X_{0}),k_{\Theta^{\prime}_{n}}(X_{\infty}))^{r}\mid X_{0},\Theta_{n},\Theta^{\prime}_{n}]]$ (16)
$=E\left[\int_{\mathcal{X}}d(k_{\Theta_{n}}(X_{0}),k_{\Theta^{\prime}_{n}}(y))^{r}\mathrm{d}\pi(y)\right]$ (17)
$=E\left[\int_{\mathcal{X}}d(k_{\Theta_{n}}(X_{0}),k_{\Theta^{\prime}_{n}}(y))^{r}\frac{\mathrm{d}\pi(y)}{\mathrm{d}\nu(y)}\mathrm{d}\nu(y)\right]$ (18)
$=E\left[E\left[d(k_{\Theta_{n}}(X_{0}),k_{\Theta^{\prime}_{n}}(Y_{0}))^{r}\frac{\mathrm{d}\pi(Y_{0})}{\mathrm{d}\nu(Y_{0})}\bigm|X_{0},\Theta_{n},\Theta^{\prime}_{n}\right]\right]$ (19)
$=E\left[d(X_{n},Y_{n})^{r}\frac{\mathrm{d}\pi(Y_{0})}{\mathrm{d}\nu(Y_{0})}\right]$ (20)
By Holder’s inequality, for $s>1$ and $K=E\left[\left(\frac{\mathrm{d}\pi(Y_{0})}{\mathrm{d}\nu(Y_{0})}\right)^{\frac{s}{s-1}}\right]^{\frac{s-1}{s}}$.
$E[d(X_{n},X_{\infty})^{r}]=E\left[d(X_{n},Y_{n})^{r}\frac{\mathrm{d}\pi(Y_{0})}{\mathrm{d}\nu(Y_{0})}\right]\leq KE\left[d(X_{n},Y_{n})^{rs}\right]^{\frac{1}{s}}$ (21)
Thus, the $L^{r}$-Wasserstein distance is bounded above as follows,
$W_{d,r}(\mathcal{L}(X_{n}),\pi)\leq E[d(X_{n},X_{\infty})^{r}]^{\frac{1}{r}}\leq K^{\frac{1}{r}}E\left[d(X_{n},Y_{n})^{rs}\right]^{\frac{1}{rs}}$ (22)
Finally by Remark 6.6 of [57], for $p\leq r$, $W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq W_{d,r}(\mathcal{L}(X_{n}),\pi)$ and so, by Equation 22,
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq W_{d,r}(\mathcal{L}(X_{n}),\pi)\leq K^{\frac{1}{r}}E\left[d(X_{n},Y_{n})^{rs}\right]^{\frac{1}{rs}}$
8.3 Alternative proof of Theorem 3.1 when $s=1$
First we define essential supremum and infimum.
Let $f:\mathcal{X}\to\mathbb{R}$ be a function and $\lambda$ be the Lebesgue measure. The essential supremum is the smallest value $a\in\mathbb{R}$ such that $\lambda(x\mid f(x)<a)=1$. More formally, $ess\sup_{x}f(x)=\inf_{a\in\mathbb{R}}\{a\mid\lambda(x\mid f(x)>a)=0\}$. The essential infimum is likewise the largest value $a\in\mathbb{R}$ such that $\lambda(x\mid f(x)>a)=1$. Or, $ess\inf_{x}f(x)=\sup_{a\in\mathbb{R}}\{a\mid\lambda(x\mid f(x)<a)=0\}$ [36].
8.3.1 Rejection rate and separation distance
Definition 8.1 (Rejection sampling).
Suppose that we have a target distribution $\pi$, which we want to sample from, but is difficult to do, and we have a proposal distribution $\nu$ that is easier to sample from. Suppose also that for the corresponding density functions $f_{\pi}(x)$ and $f_{\nu}(x)$, $f_{\nu}(x)=0\implies f_{\pi}(x)=0$, where $x\in\mathcal{X}$ and that $K=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$. To generate a random variable $X_{\infty},\mathcal{L}(X_{\infty})=\pi$ we do the following,
  1. 1.
    Sample $X\sim\nu$ and $U\sim Unif(0,1)$ independently.
  2. 2.
    If $U\leq\frac{1}{K}\frac{f_{\pi}(X)}{f_{\nu}(X)}$ then accept $X$ as a draw from $\pi$. Otherwise reject $X$ and restart from step 1.
Definition 8.2 (Rejection sampler rejection rate).
For the rejection sampler algorithm defined above, denote the event $A=\{X\text{ is accepted as a draw from }\pi\}$. The rejection rate is $r(\pi,\nu)=1-P(A)=1-1/K$ where $K=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$.
Proof.
See Section 11.2.2 of [49]. ∎
We further define the separation distance on the continuous state space $\mathcal{X}$ as follows. Separation distance was first defined in [2] for discrete state spaces. We use the definition of separation distance defined in [10] where the density functions are known.
Definition 8.3 (Separation distance (Remark 5 of [10])).
Lemma 8.1.
$s(\pi,\nu)=r(\pi,\nu)$
Proof.
$=1-\frac{1}{ess\sup_{x\in\mathcal{X}}\frac{f_{\pi}(x)}{f_{\nu}(x)}}=1-\frac{1}{K}=r(\pi,\nu)$
Alternative proof of Theorem 3.1 for when $s_{1}=1,s_{2}=\infty$.
In this proof, we use rejection sampling, which is similar to the proof method presented in [31]. Let $X_{n},Y_{n}$ be two copies of the Markov chain and $A=\{\text{Accept $Y_{0}$ as a draw from }\pi\}$. Recall from Definition 8.2 that $1-1/K=r(\pi,\nu)=1-P(A)$ and so $P(A)=1/K$. Note that $\mathcal{L}(Y_{0}|A)=\pi$ and so, $\mathcal{L}(Y_{n}|A)=\pi$.
$\geq E[d(X_{n},X_{\infty})^{p}]P(A)$
$\geq E[d(X_{n},X_{\infty})^{p}]\frac{1}{K}$
Thus for $K=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$,
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq E[d(X_{n},X_{\infty})^{p}]\leq KE[d(X_{n},Y_{n})^{p}]$
Further note that by combining Definition 8.2 and Lemma 8.1 we can write $K$ as follows
$K=\frac{1}{1-s(\pi,\nu)}=\frac{1}{1-r(\pi,\nu)}$
@@ 8.3.1 Rejection rate and separation distance
8.4 Proof of lemma related to the Gibbs sampler for a James-Stein estimator
@@
$=\frac{\Gamma(\alpha+(q-1)/2)}{\Gamma(\alpha)(2\pi\beta)^{(q-1)/2}}$